# Additive Complementary Pairs of Codes

### v2

@@ 3 Characterization of ACP of codes
Take ${\rm I\!F}_{q}={\rm I\!F}_{2}$ and ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$. Let $C$ and $D$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $3$ with generator matrices $G_{1}=\left(\begin{array}[]{ccc}1&1&0\\ \omega&0&\omega\\ 0&\omega&\omega\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{ccc}1&0&1\\ 1&1&1\\ \omega&\omega&\omega\end{array}\right),$ respectively, where $\omega$ is a primitive root of ${\rm I\!F}_{4}$. Here, we take $M$, $\pi$ and $\sigma$ are all identity. Consider a parity check matrices $H_{1}=\left(\begin{array}[]{ccc}1&1&1\\ \omega&\omega&\omega\\ \omega^{2}&\omega^{2}&0\end{array}\right)$ and $H_{2}=\left(\begin{array}[]{ccc}1&0&1\\ 1&1&0\\ \omega&0&\omega\end{array}\right)$ of $C$ and $D$, respectively. Then $H_{2}G_{1}^{\top}=\left(\begin{array}[]{ccc}1&0&\omega\\ 1&\omega&\omega\\ \omega&0&\omega^{2}\end{array}\right)$ and $H_{1}G_{2}^{\top}=\left(\begin{array}[]{ccc}0&1&\omega\\ 0&\omega&\omega^{2}\\ \omega^{2}&0&0\end{array}\right)$. Here, $Tr(H_{1}G_{2}^{\top})=\left(\begin{array}[]{ccc}0&0&1\\ 0&1&1\\ 1&0&0\end{array}\right)$ and $Tr(H_{2}G_{1}^{\top})=\left(\begin{array}[]{ccc}0&0&1\\ 0&1&1\\ 1&0&1\end{array}\right)$ which are having rank $3$. Hence, $(C,D)$ is ACP over ${\rm I\!F}_{4}$ (see Theorem 3.6).
@@ 4 Building-up construction for ACP of codes
It can be shown that $G$ is a generator matrix of an ${\rm I\!F}_{q}$-linear additive code, say $C$, corresponding to the linear code $\tilde{C}$ over ${\rm I\!F}_{q^{m}}$. Next, we will present a construction of ACP of codes from LCP of codes over ${\rm I\!F}_{q^{m}}$. Towards the LCP of codes, we define a general inner product on ${\rm I\!F}_{q^{m}}$, which is the analog of the inner product $\mathcal{B}$ (see Equation (2.1)), using the bilinear mapping
@@
where $\mu_{i},\sigma$, and $\pi$ are as defined in Equation (2.1). Notice that, in general, $\tilde{\mathcal{B}}$ is not symmetric. A pair of codes $(\tilde{C},\tilde{D})$ of length $n$ over ${\rm I\!F}_{q^{m}}$ is called an LCP of codes if $\tilde{C}\oplus_{{\rm I\!F}_{q^{m}}}\tilde{D}={\rm I\!F}_{q^{m}}^{n}$. That is, a pair of codes $(\tilde{C},\tilde{D})$ of length $n$ over ${\rm I\!F}_{q^{m}}$ is LCP if and only if $\tilde{C}\cap\tilde{D}=\{0\}$ and $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{D})=n$. The following proposition presents a characterization of LCP of codes over ${\rm I\!F}_{q^{m}}$ with respect to the inner product $\tilde{\mathcal{B}}$ (see Equation (4)).
@@
Let $\tilde{C}$ and $\tilde{D}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$ with generator matrices $\tilde{G}_{1}$, $\tilde{G}_{2}$ and parity check matrices $\tilde{H}_{1}$, $\tilde{H}_{2}$, respectively. Then the followings are equivalent.
@@
    The pair $(\tilde{C},\tilde{D})$ is LCP.
@@
    The $rank\left(\left(\tilde{H}_{2}D\pi\left(\tilde{G}_{1}\right)P\right)^{\top}\right)=rank(\tilde{G}_{1})$ and $rank\left(Tr\left(\tilde{G}_{1}D\pi\left(\tilde{G}_{2}\right)P\right)^{\top}\right)=rank(\tilde{G}_{2})$. Additionally, $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{D})=n$.
@@
The proof of this Proposition is similar to the proof of the Theorem 3.6. Now we will present an ACP of codes from a given pair of LCP of codes.
@@
Let $\tilde{C}$, $\tilde{D}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$. Let $C$, $D$ be the codes constructed from $\tilde{C}$, $\tilde{D}$, respectively, as presented in Equation (13). If the pair $(\tilde{C},\tilde{D})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$ then $(C,D)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$.
@@
Let $\tilde{G}_{1}$, $\tilde{G}_{2}$ be generator matrices of the linear codes $\tilde{C}$, $\tilde{D}$, respectively. Using Equation 13, construct $G_{1}$ and $G_{2}$ which are generator matrices of $C$, $D$, respectively. Since $(\tilde{C},\tilde{D})$ is an LCP of codes, $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{D})=n$ and that implies $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=mn$. We need to prove that $C\cap D=\{0\}$. In contrary, assume that there exists a non-zero $x\in{\rm I\!F}_{q}^{n}$ such that $x\in C\cap D$. Then $x=\gamma G_{1}=\delta G_{2}$ for some $\gamma\in{\rm I\!F}_{q}^{k}$, $\delta\in{\rm I\!F}_{q}^{mn-k}$ where $k=rank(G_{1})$ and $mn-k=rank(G_{2})$. Then from Equation  (13), we have that $x\in\tilde{C}\cap\tilde{D}$. Since $x$ is nonzero vector in ${\rm I\!F}_{q}^{n}$, it contradicts that $(\tilde{C},\tilde{D})$ is an LCP. Hence, $(C,D)$ is an ACP of codes. ∎
Converse of Proposition 4.2 is not correct in general. For this, we demonstrate an example as follows.
Example 4.1.
Take ${\rm I\!F}_{q}={\rm I\!F}_{2}$ and ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$. Let $C$ and $D$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $3$ with generator matrices $G_{1}=\left(\begin{array}[]{ccc}1&1&0\\ \omega&0&\omega\\ 0&\omega&\omega\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{ccc}1&0&1\\ 1&1&1\\ \omega&\omega&\omega\end{array}\right),$ respectively, where $\omega$ is a primitive root of ${\rm I\!F}_{4}$. Here, $(C,D)$ is ACP over ${\rm I\!F}_{4}$. It is easy to see that $\tilde{C}\cap\tilde{D}\neq\{0\}$, as $(\omega~{}0~{}\omega)\in\tilde{C}\cap\tilde{D}$. Therefore, $(\tilde{C},\tilde{D})$ is not an LCP of codes over ${\rm I\!F}_{4}$.
For $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})\in\left({\rm I\!F}_{q^{m}}^{n}\right)$ and a code $C$, define
${\mathbf{a}}C:=\{(a_{1}c_{1},a_{2}c_{2},\ldots,a_{n}c_{n})~{}|~{}(c_{1},c_{2},\ldots,c_{n}\in C)\}.$
Note that, for $\mathbf{a}\in\left({\rm I\!F}_{q^{m}}^{*}\right)^{n}$, ${\mathbf{a}}C$ is an ${\rm I\!F}_{q}$-linear additive code if $C$ is an ${\rm I\!F}_{q}$-linear additive code. Recall that a linear code $\tilde{C}:=[n,k,d]$ over ${\rm I\!F}_{q^{m}}$ is MDS if $d=n-k+1$.
@@
Let $\tilde{C}$, $\tilde{D}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$. Let $C$, $D$ be the codes constructed from $\tilde{C}$, $\tilde{D}$, respectively, as presented in Equation (13). Assume that at least one of codes $\tilde{C},\tilde{D}$ is MDS and $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{D})=n$. Then there exists $\mathbf{a}\in\left({\rm I\!F}_{q^{m}}^{*}\right)^{n}$ such that a pair $({\mathbf{a}}C,D)$ is an ACP of codes.
@@
Since $q^{m}\geq 4$, by [1, Theorm 5.4], there exists $\mathbf{a}\in\left({\rm I\!F}_{q^{m}}^{*}\right)^{n}$ such that $(\mathbf{a}\tilde{C},\tilde{D})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$. By Proposition 4.2, we obtain that $({\mathbf{a}}C,D)$ is an ACP of codes. ∎
@@
Let $\tilde{C}$ be a linear code over ${\rm I\!F}_{q^{m}}$ of length $n$ with generator matrix $\tilde{G}$. We define two different expansions of $\tilde{C}$ as $\tilde{C}_{ex_{1}}$ and $\tilde{C}_{ex_{2}}$ generated by the generator matrices
@@
Let $\tilde{C}$ and $\tilde{D}$ be two linear codes over ${\rm I\!F}_{q^{m}}$ such that a pair $(\tilde{C},\tilde{D})$ is an ACP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1). Then there exist $\lambda\in{\rm I\!F}_{q^{m}}^{*}$ and $P\in\left({\rm I\!F}_{q^{m}}^{n}\right)^{*}$ such that the pair $(\tilde{C}_{ex_{1}},\tilde{D}_{ex_{2}})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$.
@@
Let $\tilde{C}$ and $\tilde{D}$ be two linear codes over ${\rm I\!F}_{q^{m}}$ with generator matrices $\tilde{G}_{1}$ and $\tilde{G}_{2}$, respectively. Consider the codes $\tilde{C}_{ex_{1}}$ and $\tilde{D}_{ex_{2}}$ with generator matrices $\tilde{G}_{ex_{1}}=\left(\begin{array}[]{ccc}\lambda&P\\ 0&\tilde{G}_{1}\end{array}\right)$ and $\tilde{G}_{ex_{2}}=\left(\begin{array}[]{ccc}P^{\prime\top}&\tilde{G}_{2}\end{array}\right)$, respectively, where $\lambda\in{\rm I\!F}_{q^{m}}$, non-zero $P$ is an $(1\times n)$ matrix over ${\rm I\!F}_{q^{m}}$ and $P^{\prime}$ is an $1\times(n-k)$ sub-matrix of $P$. Since, a pair $(\tilde{C},\tilde{D})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$, then by Proposition 4.1, $rank\left(\begin{array}[]{ccc}\tilde{G}_{1}\\ \tilde{G}_{2}\end{array}\right)=n$. Hence,

### v3

@@ 3 Characterization of ACP of codes
Take ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$ with a primitive element $\omega$. Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $3$ with generator matrices $G_{1}=\left(\begin{array}[]{ccc}1&1&0\\ \omega&0&\omega\\ 0&\omega&\omega\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{ccc}1&0&1\\ 1&1&1\\ \omega&\omega&\omega\end{array}\right)$, respectively. Here, we take $M,\pi$ and $\sigma$ are all identity.
Consider a parity check matrices $H_{1}=\left(\begin{array}[]{ccc}1&1&1\\ \omega&\omega&\omega\\ \omega^{2}&\omega^{2}&0\end{array}\right)$ and $H_{2}=\left(\begin{array}[]{ccc}1&0&1\\ 1&1&0\\ \omega&0&\omega\end{array}\right)$ of $\mathcal{C}$ and $\mathcal{D}$, respectively. Then $H_{1}G_{2}^{\top}=\left(\begin{array}[]{ccc}0&1&\omega\\ 0&\omega&\omega^{2}\\ \omega^{2}&0&0\end{array}\right)$ and $H_{2}G_{1}^{\top}=\left(\begin{array}[]{ccc}1&0&\omega\\ 1&\omega&\omega\\ \omega&0&\omega^{2}\end{array}\right)$. Here, $Tr(H_{1}G_{2}^{\top})=\left(\begin{array}[]{ccc}0&0&1\\ 0&1&1\\ 1&0&0\end{array}\right)$ and $Tr(H_{2}G_{1}^{\top})=\left(\begin{array}[]{ccc}0&0&1\\ 0&1&1\\ 1&0&1\end{array}\right)$ which are having rank $3$. Hence, $(\mathcal{C},\mathcal{D})$ is ACP over ${\rm I\!F}_{4}$ (see Theorem 3.7).
Example 3.9.
Take ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$ with a primitive element $\omega$ and an automorphism $\pi$ on ${\rm I\!F}_{q^{m}}$ defined as $\pi(x)=x^{2}$. Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $6$ with generator matrices
$G_{1}=\left(\begin{array}[]{cccccc}1&0&0&0&0&1\\ \omega&0&0&0&0&\omega\\ 0&1&0&0&0&1\\ 0&\omega&0&0&0&\omega\\ 0&0&1&0&0&1\\ 0&0&\omega&0&0&\omega\\ 0&0&0&1&0&1\\ 0&0&0&\omega&0&\omega\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{cccccc}0&0&0&0&1&0\\ 0&0&0&0&\omega&0\\ 1&1&1&1&0&1\\ \omega&\omega&\omega&\omega&0&\omega\end{array}\right)$, respectively.
It is easy to see that $(\mathcal{C},\mathcal{D})$ is an ACP of codes with respect to $P$ such that
$P_{ij}=\left\{\begin{array}[]{ll}1&\textrm{if~{} off-diagonal element };\\ 0&\textrm{otherwise,}\end{array}\right.$
$M=diag(\mu_{1},\mu_{2},\ldots,\mu_{n})$ is an $n\times n$ matrix with $\mu_{i}\in{\rm I\!F}_{4}^{*}$. However, $rank\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ \end{array}\right)\right)=6$, which coincides with Theorem 3.7, since $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1})$ and
$rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$.
@@ 4 Building-up construction for ACP of codes
It can be shown that $G$ is a generator matrix of an ${\rm I\!F}_{q}$-linear additive code, say $\mathcal{C}$, corresponding to the linear code $\tilde{\mathcal{C}}$ over ${\rm I\!F}_{q^{m}}$, which is nothing but $\tilde{\mathcal{C}}$, i.e., $\mathcal{C}=\tilde{\mathcal{C}}$. Next, we will present a construction of ACP of codes from LCP of codes over ${\rm I\!F}_{q^{m}}$. Towards the LCP of codes, we define a general inner product on ${\rm I\!F}_{q^{m}}$, which is the analogue of the inner product $\mathcal{B}$ (see Equation (2.1)), using the bilinear mapping
@@
where $\mu_{i},\sigma$, and $\pi$ are as defined in Equation (2.1). Notice that, in general, $\tilde{\mathcal{B}}$ is not symmetric. A pair of codes $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ of length $n$ over ${\rm I\!F}_{q^{m}}$ is called an LCP of codes if $\tilde{\mathcal{C}}\oplus_{{\rm I\!F}_{q^{m}}}\tilde{\mathcal{D}}={\rm I\!F}_{q^{m}}^{n}$. That is, a pair of codes $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ of length $n$ over ${\rm I\!F}_{q^{m}}$ is LCP if and only if $\tilde{\mathcal{C}}\cap\tilde{\mathcal{D}}=\{0\}$ and $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{D}})=n$. The following proposition presents a characterization of LCP of codes over ${\rm I\!F}_{q^{m}}$ with respect to the inner product $\tilde{\mathcal{B}}$ (see Equation (4)).
@@
Let $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{D}}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$ with generator matrices $\tilde{G}_{1}$, $\tilde{G}_{2}$ and parity check matrices $\tilde{H}_{1}$, $\tilde{H}_{2}$, respectively. Then, the following are equivalent.
@@
    The pair $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is LCP.
@@
    The $rank\left(\left(\tilde{H}_{2}M\left(\pi\left(\tilde{G}_{1}\right)P\right)^{\top}\right)\right)=rank(\tilde{G}_{1})$ and $rank\left(Tr\left(\tilde{G}_{1}M\left(\pi\left(\tilde{G}_{2}\right)P\right)^{\top}\right)\right)=rank(\tilde{G}_{2})$. Additionally, $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{D}})=n$.
@@
The proof of this Proposition is similar to the proof of the Theorem 3.7. Now, we will present an ACP of codes from a given pair of LCP of codes.
@@
Let $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$. Let $\mathcal{C}$, $\mathcal{D}$ be the codes constructed from $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$, respectively, as presented in Equation (11). If the pair $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$ then $(\mathcal{C},\mathcal{D})$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$.
@@
Let $\tilde{G}_{1}$, $\tilde{G}_{2}$ be generator matrices of the linear codes $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$, respectively. Using Equation 11, construct $G_{1}$ and $G_{2}$ which are generator matrices of $\mathcal{C}$, $\mathcal{D}$, respectively. Since $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes, $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{D}})=n$ and that implies $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=mn$. We need to prove that $\mathcal{C}\cap\mathcal{D}=\{0\}$. On the contrary, assume that there exists a non-zero $\mathbf{x}\in{\rm I\!F}_{q}^{n}$ such that $\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$. Then $\mathbf{x}=\gamma G_{1}=\delta G_{2}$ for some $\gamma\in{\rm I\!F}_{q}^{k}$, $\delta\in{\rm I\!F}_{q}^{mn-k}$ where $k=rank(G_{1})$ and $mn-k=rank(G_{2})$. Then from Equation  (11), we have that $\mathbf{x}\in\tilde{\mathcal{C}}\cap\tilde{\mathcal{D}}$. Since $\mathbf{x}$ is nonzero vector in ${\rm I\!F}_{q}^{n}$, it contradicts that $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP. Hence, $(\mathcal{C},\mathcal{D})$ is an ACP of codes. ∎
Towards this, for an ${\rm I\!F}_{q}$-linear additive code $C$ over ${\rm I\!F}_{q^{m}}$ of length $n$, we define
$Tr(\mathcal{C}):=\{Tr(\mathbf{c}):\mathbf{c}\in\mathcal{C}\}.$
Note that $Tr(\mathcal{C})$ is a linear code over ${\rm I\!F}_{q}$ of length $n$. We call $Tr(\mathcal{C})$, a trace code of $\mathcal{C}$. Further, if $G$ is a generator matrix of $\mathcal{C}$, then $Tr(G)$ (defined in Section 3) is a generator matrix of $Tr(\mathcal{C})$.
@@
Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$. If the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then $(Tr(\mathcal{C}),Tr(\mathcal{D}))$ is an LCP of codes over ${\rm I\!F}_{q}$. In particular, if $\mathcal{C}$ is an additive complementary dual (ACD) codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then $Tr(\mathcal{C})$ is an LCD codes over ${\rm I\!F}_{q}$.
@@
Let $G_{1}$ and $G_{2}$ be two generator matrices of $\mathcal{C}$ and $\mathcal{D}$, respectively. Since, the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then by Theorem 3.3, we obtain that $rank\left(Tr\left(\begin{array}[]{cc}G_{1}\\ G_{2}\end{array}\right)\right)=n$. That is, $rank\left(\begin{array}[]{cc}Tr(G_{1})\\ Tr(G_{2})\end{array}\right)=n$. Since $Tr(G_{1})$ and $Tr(G_{2})$ are generator matrices of $Tr(\mathcal{C})$ and $Tr(\mathcal{D})$, respectively, from Proposition 4.1, we get $(Tr(\mathcal{C}),Tr(\mathcal{D}))$ is an LCP of codes over ${\rm I\!F}_{q}$. ∎
Theorem 4.2.
Let $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$. Let $\mathcal{C}$, $\mathcal{D}$ be the codes constructed from $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$, respectively, as presented in Equation (11). Then the pair $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$ if and only if $(\mathcal{C},\mathcal{D})$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1).
Proof.
Combining Proposition 4.2 and Theorem 4.1, we have the required result. ∎
For $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})\in{\rm I\!F}_{q^{m}}^{n}$ and a code $\mathcal{C}$, define
${\mathbf{a}}\mathcal{C}:=\{(a_{1}c_{1},a_{2}c_{2},\ldots,a_{n}c_{n})~{}|~{}(c_{1},c_{2},\ldots,c_{n})\in\mathcal{C})\}.$
Note that, for $\mathbf{a}\in\left({\rm I\!F}_{q^{m}}^{*}\right)^{n}$, ${\mathbf{a}}\mathcal{C}$ is an ${\rm I\!F}_{q}$-linear additive code if $C$ is an ${\rm I\!F}_{q}$-linear additive code. Recall that a linear code $\tilde{\mathcal{C}}:=[n,k,d]$ over ${\rm I\!F}_{q^{m}}$ is MDS if $d=n-k+1$.
Theorem 4.3.
Let $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$ be two linear codes of length $n$ over ${\rm I\!F}_{q^{m}}$. Let $\mathcal{C}$, $\mathcal{D}$ be the codes constructed from $\tilde{\mathcal{C}}$, $\tilde{\mathcal{D}}$, respectively, as presented in Equation (11). Assume that at least one of codes $\tilde{\mathcal{C}},\tilde{\mathcal{D}}$ is MDS and $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{D}})=n$. Then there exists $\mathbf{a}\in\left({\rm I\!F}_{q^{m}}^{*}\right)^{n}$ such that a pair $({\mathbf{a}}\mathcal{C},\mathcal{D})$ is an ACP of codes.
Proof.
Since $q^{m}\geq 4$, by [1, Theorem 5.4], there exists $\mathbf{a}\in\left({\rm I\!F}_{q^{m}}^{*}\right)^{n}$ such that $(\mathbf{a}\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$. By Proposition 4.2, we obtain that $({\mathbf{a}}\mathcal{C},\mathcal{D})$ is an ACP of codes. ∎
@@
Now, we will present a numerical example.
Example 4.5.
Let ${\rm I\!F}_{25}$ be a finite field, $x$ be transcendental over ${\rm I\!F}_{25}$ and $\mathcal{P}_{2}=\{f\in{\rm I\!F}_{q}[x]:\deg(f)\leq 1\}$. For $\textbf{b}=(1,2,3,4)\in{\rm I\!F}_{25}^{4}$, let $RS_{2}(\textbf{b})=\{(f(1),f(2),f(3),f(4))~{}|~{}f\in\mathcal{P}_{2}\}$ be a $2$-dimensional Reed-Solomon code which is MDS. Then $RS_{2}(\textbf{b})$ is $[4,2,3]$ code and $RS_{2}(\textbf{b})^{\perp}$ is a $[4,2,3]$ code with generator matrix $G=\left(\begin{array}[]{cccc}1&1&1&1\\ 1&2&3&4\end{array}\right)$ and parity check matrix $H=\left(\begin{array}[]{cccc}1&0&2&2\\ 0&1&3&1\end{array}\right)$. Set $\mathbf{a}=(2,1,1,1)$, then $(\mathbf{a}RS_{2}(\textbf{b}),RS_{2}(\textbf{b})^{\perp})$ forms an LCP over ${\rm I\!F}_{25}$. It is easy to see that $({\mathbf{a}}\widehat{RS_{2}(\textbf{b})},\widehat{RS_{2}(\textbf{b})^{\perp}})$ is an ACP of codes over ${\rm I\!F}_{25}$.
Let $\tilde{C}$ be a linear code over ${\rm I\!F}_{q^{m}}$ of length $n$ with generator matrix $\tilde{G}$. We define two different expansions of $\tilde{\mathcal{C}}$ as $\tilde{\mathcal{C}}_{ex_{1}}$ and $\tilde{\mathcal{C}}_{ex_{2}}$ generated by the generator matrices
@@
Let $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{D}}$ be two linear codes over ${\rm I\!F}_{q^{m}}$ such that a pair $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1). Then there exist $\lambda\in{\rm I\!F}_{q^{m}}^{*}$ and $P\in{\rm I\!F}_{q^{m}}^{n}$ such that the pair $(\tilde{\mathcal{C}}_{ex_{1}},\tilde{\mathcal{D}}_{ex_{2}})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$.
@@
Let $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{D}}$ be two linear codes over ${\rm I\!F}_{q^{m}}$ with generator matrices $\tilde{G}_{1}$ and $\tilde{G}_{2}$, respectively. Consider the codes $\tilde{\mathcal{C}}_{ex_{1}}$ and $\tilde{\mathcal{D}}_{ex_{2}}$ with generator matrices $\tilde{G}_{ex_{1}}=\left(\begin{array}[]{ccc}\lambda&P\\ 0&\tilde{G}_{1}\end{array}\right)$ and $\tilde{G}_{ex_{2}}=\left(\begin{array}[]{ccc}P^{\prime\top}&\tilde{G}_{2}\end{array}\right)$, respectively, where $\lambda\in{\rm I\!F}_{q^{m}}$, non-zero $P$ is an $(1\times n)$ matrix over ${\rm I\!F}_{q^{m}}$ and $P^{\prime}$ is an $1\times(n-k)$ sub-matrix of $P$. Since, a pair $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$, then by Proposition 4.1, $rank\left(\begin{array}[]{ccc}\tilde{G}_{1}\\ \tilde{G}_{2}\end{array}\right)=n$. Hence,